OpenStudy (anonymous):

A design on the surface of a balloon is 4 cm wide when the balloon holds 53 cm3 of air. How much air does the balloon hold when the design is 16 cm wide?

jimthompson5910 (jim_thompson5910):

Hint: solve for x (4/16)^3 = 53/x

OpenStudy (anonymous):

how do I do that

jimthompson5910 (jim_thompson5910):

(4/16)^3 = 53/x (1/4)^3 = 53/x 1/64 = 53/x 1*x = 64*53 x = ______

OpenStudy (anonymous):

3392?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

yay!

jimthompson5910 (jim_thompson5910):

this is all based on the fact that all spheres are similar solids so we can set up this ratio and solve